# ConCap: Practical Network Traffic Generation for Flow-based Intrusion Detection Systems

### v1

@@ A-A NetFlow Exporter Configuration
containerImage: ghcr.io/anonymous/concap/argus:latest
@@
A-B Configuration of a DetGen Scenario in ConCap
To demonstrate the flexibility of ConCap, we have created a step-by-step guide to implement one of the predefined scenarios from DetGen [22]. We selected the well-documented “capture-020-nginx” scenario, which uses “siege”, an HTTP load testing and benchmarking tool, to target the “nginx” HTTP server and reverse proxy. In this guide, we show how to replicate the “capture-020-nginx” scenario in ConCap using processing pods and scenario definitions.
@@ A-B Configuration of a DetGen Scenario in ConCap
image: ghcr.io/anonymous/concap/siege:ubuntu18
@@
image: ghcr.io/anonymous/concap/advanced-attacker:1.0.0
@@
image: ghcr.io/anonymous/concap/openssh-server:password-24.04
@@
image: ghcr.io/anonymous/concap/mysql:1.0.0
@@
A-C Scenario Details for Broad Analysis ConCap
@@
Patator FTP This scenario emulates an FTP brute-force attack using Patator. A password list with the 200 most used passwords of 2023 [109] was explicitly supplied, and failure responses matching the message ”530 Login incorrect.” were ignored using Patator’s ignore option.
@@
Patator SSH This scenario represents a brute-force attack against an SSH server. Both a username and password list were provided, top usernames shortlist and top 200 passwords from 2023 [109], and the failed login attempts are filtered based on the message ”Authentication failed.”
@@
Appendix B Datasets (existing and new ones)
@@
B-A Real-word Network Capture: Description
@@
B-B Preprocessing of Benchmark datasets
@@
NetFlow Extraction The fixed versions of the datasets CICIDS17 and CICIDS18 were released in October 2022. Since then, CICFlowMeter, the tool used to extract the NetFlow features from the network traffic traces, has received over 30 new commits fixing, changing, and adding NetFlow features. To benefit from these updates over the last two years, we replicated the work done by Liu et al. [6] by extracting the NetFlows using the current version (Aug ’24) of CICFlowMeter and labeling the flows based on their fixed logic. The “attempted” NetFlows are removed from our dataset.
Data Cleaning. The dataset cleaning steps on the NetFlows performed in our experiments follow the best practices described in previous work [110, 4, 111]. First, the meta-data and spurious features are removed: “id”, “Flow ID”, “Src IP”, “Dst IP”, “Timestamp”, “FWD Init Win Bytes”, and “Bwd Init Win Bytes”. Second, the source and destination ports are mapped to their IANA port categories and subsequently encoded as 0, 1, or 2 for respectively well-known, registered, and dynamic. Third, all NetFlows with missing values are removed. Last, all duplicate NetFlows are removed.
@@
B-C Generation of (new) CVE Data with ConCap
We generated new labeled ”malicious” data using ConCap by executing three recent CVE-based attack scenarios, ensuring none of them overlap with the datasets mentioned in Flood et al. [8]. These CVEs represent diverse vulnerabilities, each exploited to demonstrate the flexibility of ConCap in handling modern attacks.
  • CVE-2024-47177 [77]: This vulnerability affects OpenPrinting Cups-Browsed versions 2.0.1 and earlier, where improper handling of the FoomaticRIPCommandLine parameter in PostScript Printer Description (PPD) files allows remote code execution. Attackers can exploit this by creating a malicious IPP (Internet Printing Protocol) server that sends crafted printer information to a vulnerable Cups-Browsed instance, enabling arbitrary command execution on the affected system.
  • CVE-2024-36401 [78]: In GeoServer versions prior to 2.25.1, 2.24.3, and 2.23.5, unauthenticated remote code execution is possible through unsafely evaluated property name expressions. Specially crafted input can exploit these unsafe evaluations in multiple OGC request parameters, allowing attackers to execute arbitrary code on a vulnerable GeoServer installation.
  • CVE-2024-2961 [79]: This vulnerability in the GNU C Library (versions 2.39 and earlier) affects the iconv() function, which may overflow the output buffer by up to 4 bytes when converting strings to the ISO-2022-CN-EXT character set. Attackers can exploit arbitrary file read vulnerabilities in PHP applications to escalate to remote code execution by leveraging the iconv() issue, potentially crashing the application or executing code.
@@ B-C Generation of (new) CVE Data with ConCap
Appendix C Additional Results
@@
We provide the tables with the standard deviation of our results (useful to carry out statistical tests) in our repository [23].
C-A Runtime and Computing Results
We report in Table VII the training time of the ML models used for the experiments in §VI. These experiments have been carried out on a machine with a CPU and RAM.
@@ C-A Runtime and Computing Results
Baseline Adversarial Open World
Train Set Benign + P=1 Benign + P=0 Benign + P=1 + P=0 Benign + P=0 + P=1 Benign + P=0 + P=1
DT $0.3\pm 0.0$.0) $0.2\pm 0.0$.0) $0.4\pm 0.0$.0) $0.5\pm 0.0$.1) $0.4\pm 0.0$.0)
RF $3.1\pm 0.0$.1) $4.1\pm 0.0$.2) $5.4\pm 0.0$.2) $5.4\pm 0.0$.2) $5.4\pm 0.0$.1)
XGB $3.0\pm 0.0$.4) $3.9\pm 0.0$.6) $4.5\pm 0.0$.3) $3.9\pm 0.0$.2) $5.1\pm 0.0$.6)
SVM $0.5\pm 0.0$.0) $0.9\pm 0.0$.1) $1.2\pm 0.0$.2) $1.1\pm 0.0$.0) $1.1\pm 0.0$.1)
DNN $9.2\pm 0.0$.5) $12.3\pm 0.0$.5) $13.9\pm 0.0$.8) $11.7\pm 0.0$.2) $11.8\pm 0.0$.8)
@@
Baseline Adversarial Open World
DT $6.7\pm 0.1$.3) $8.2\pm 0.1$.2) $5.3\pm 0.0$.6)
RF $40.8\pm 0.4$.2) $36.9\pm 0.1$.4) $38.0\pm 0.3$.6)
XGB $15.1\pm 0.3$.8) $21.9\pm 0.3$.4) $27.1\pm 0.6$.6)
SVM $4.1\pm 0.0$.9) $12.5\pm 0.1$.4) $25.3\pm 0.6$.5)
DNN $17.2\pm 0.2$.0) $22.6\pm 0.0$.5) $20.0\pm 0.1$.5)
@@
Baseline Adversarial Open World
DT $40.4\pm 0.1$.8) $38.6\pm 0.3$.3) $35.1\pm 0.3$.2)
RF $547.8\pm 0.6$.3) $618.5\pm 5.5$.9) $504.3\pm 4.1$.9)
XGB $118.8\pm 0.7$.4) $120.8\pm 0.7$.6) $47.8\pm 3.0$.8)
SVM $101.3\pm 1.0$.9) $131.1\pm 1.8$.3) $389.6\pm 5.8$.2)
DNN $574.5\pm 4.5$.8) $384.3\pm 2.8$.4) $389.3\pm 4.6$.6)
@@
TABLE VIII: Performance evaluation of ConCap running on a multi-node K8s cluster versus microK8s on a WSL-enabled laptop.

### v2

@@ B-A NetFlow Exporter Configuration
containerImage: ghcr.io/idlab-discover/concap/argus:latest
@@
B-B Configuration of a DetGen Scenario in ConCap
To demonstrate the flexibility of ConCap, we have created a step-by-step guide to implement one of the predefined scenarios from DetGen clausen2019traffic . We selected the well-documented “capture-020-nginx” scenario, which uses “siege”, an HTTP load testing and benchmarking tool, to target the “nginx” HTTP server and reverse proxy. In this guide, we show how to replicate the “capture-020-nginx” scenario in ConCap using processing pods and scenario definitions.
@@ B-B Configuration of a DetGen Scenario in ConCap
image: ghcr.io/idlab-discover/concap/siege:ubuntu18
@@
image: ghcr.io/idlab-discover/concap/advanced-attacker:1.0.0
@@
image: ghcr.io/idlab-discover/concap/openssh-server:password-24.04
@@
image: ghcr.io/idlab-discover/concap/mysql:1.0.0
@@
B-C Configuration of a SOCBED Scenario in ConCap
We further show the coverage of ConCap by providing the configuration files that enable reproduction of the experimental setup of SOCBED uetz2021reproducible .
To this end, we report in Listing 6 the configuration that reproduces a portscan, i.e., the first step of the attack chain in SOCBED. We also provide the corresponding PCAP and (labeled) NetFlow data in our repository (note: this data is not provided by SOCBED). Furthermore, in our repository repository , we report additional configuration files (and corresponding data) for other scenarios in SOCBED.
We observe that creating all these configuration files required us less than one day of work: we merely investigated the public repository of SOCBED socbedGithub to infer the low-level network and command details, required to devise our own configuration files.
Listing 6: A configuration of ConCap to support one of the scenarios envisioned in SOCBED which includes a portscan.
type: multi-target
name: socbed-example
attacker:
name: attacker
image: instrumentisto/nmap
atkCommand: nmap \$TARGET_IP_1 \$TARGET_IP_2 \$TARGET_IP_3 \$TARGET_IP_4 \$TARGET_IP_5 -n disable-arp-ping -sU -sV
cpuRequest: 200m
memRequest: 200Mi
targets:
- name: dmz-server
image: vulnerables/web-dvwa
cpuRequest: 200m
memRequest: 200Mi
labels:
server: dmz-server
- name: log-server
image: kibana:9.2.1
cpuRequest: 200m
memRequest: 200Mi
labels:
server: log-server
- name: internal-server
image: dperson/samba
cpuRequest: 200m
memRequest: 200Mi
labels:
server: internal-server
- name: client-1
image: dockurr/windows
cpuRequest: 200m
memRequest: 200Mi
labels:
server: client-1”
- name: client-2
image: dockurr/windows
cpuRequest: 200m
memRequest: 200Mi
labels:
server: client-2”
network:
bandwidth: 100Mbit
queueSize: 100ms
delay: 5ms
labels:
label: 1
category: socbed-example
B-D Scenario Details for Broad Analysis ConCap
@@
Patator FTP This scenario emulates an FTP brute-force attack using Patator. A password list with the 200 most used passwords of 2023 danielmiesslerseclists was explicitly supplied, and failure responses matching the message ”530 Login incorrect.” were ignored using Patator’s ignore option.
@@
Patator SSH This scenario represents a brute-force attack against an SSH server. Both a username and password list were provided, top usernames shortlist and top 200 passwords from 2023 danielmiesslerseclists , and the failed login attempts are filtered based on the message ”Authentication failed.”
@@
Appendix C Datasets (existing and new ones)
@@
C-A Real-word Network Capture: Description
@@
C-B Preprocessing of Benchmark datasets
@@
NetFlow Extraction The fixed versions of the datasets CICIDS17 and CICIDS18 were released in October 2022. Since then, CICFlowMeter, the tool used to extract the NetFlow features from the network traffic traces, has received over 30 new commits fixing, changing, and adding NetFlow features. To benefit from these updates over the last two years, we replicated the work done by Liu et al. liu2022error by extracting the NetFlows using the current version (Aug ’24) of CICFlowMeter and labeling the flows based on their fixed logic. The “attempted” NetFlows are removed from our dataset.
Data Cleaning. The dataset cleaning steps on the NetFlows performed in our experiments follow the best practices described in previous work apruzzese2022sok ; arp2022and ; d2022establishing . First, the meta-data and spurious features are removed: “id”, “Flow ID”, “Src IP”, “Dst IP”, “Timestamp”, “FWD Init Win Bytes”, and “Bwd Init Win Bytes”. Second, the source and destination ports are mapped to their IANA port categories and subsequently encoded as 0, 1, or 2 for respectively well-known, registered, and dynamic. Third, all NetFlows with missing values are removed. Last, all duplicate NetFlows are removed.
@@
C-C Generation of (new) CVE Data with ConCap
We generated new labeled ”malicious” data using ConCap by executing three recent CVE-based attack scenarios, ensuring none of them overlap with the datasets mentioned in Flood et al. flood2024bad . These CVEs represent diverse vulnerabilities, each exploited to demonstrate the flexibility of ConCap in handling modern attacks.
  • CVE-2024-47177 cve1 : This vulnerability affects OpenPrinting Cups-Browsed versions 2.0.1 and earlier, where improper handling of the FoomaticRIPCommandLine parameter in PostScript Printer Description (PPD) files allows remote code execution. Attackers can exploit this by creating a malicious IPP (Internet Printing Protocol) server that sends crafted printer information to a vulnerable Cups-Browsed instance, enabling arbitrary command execution on the affected system.
  • CVE-2024-36401 cve2 : In GeoServer versions prior to 2.25.1, 2.24.3, and 2.23.5, unauthenticated remote code execution is possible through unsafely evaluated property name expressions. Specially crafted input can exploit these unsafe evaluations in multiple OGC request parameters, allowing attackers to execute arbitrary code on a vulnerable GeoServer installation.
  • CVE-2024-2961 cve3 : This vulnerability in the GNU C Library (versions 2.39 and earlier) affects the iconv() function, which may overflow the output buffer by up to 4 bytes when converting strings to the ISO-2022-CN-EXT character set. Attackers can exploit arbitrary file read vulnerabilities in PHP applications to escalate to remote code execution by leveraging the iconv() issue, potentially crashing the application or executing code.
@@ C-C Generation of (new) CVE Data with ConCap
Appendix D Additional Results and Experiments
@@
We provide the tables with the standard deviation of our results (useful to carry out statistical tests) in our repository repository .
D-A Runtime and Computing Results
We report in Table VII the training time of the ML models used for the experiments in §VI. These experiments have been carried out on a machine running Windows 11 Pro (Build 26100) with a Intel 12th Gen Core i7-1265U CPU 1.80 GHz (10 cores, 12 threads) and 32GB RAM.
@@ D-A Runtime and Computing Results
Train Set Benign + P=1 Benign + P=0
DT $0.3\pm 0.0$.0) $0.2\pm 0.0$.0)
RF $3.1\pm 0.0$.1) $4.1\pm 0.0$.2)
XGB $3.0\pm 0.0$.4) $3.9\pm 0.0$.6)
SVM $0.5\pm 0.0$.0) $0.9\pm 0.0$.1)
DNN $9.2\pm 0.0$.5) $12.3\pm 0.0$.5)
@@
Train time
DT $6.7\pm 0.1$.3)
RF $40.8\pm 0.4$.2)
XGB $15.1\pm 0.3$.8)
SVM $4.1\pm 0.0$.9)
DNN $17.2\pm 0.2$.0)
@@
Train time
DT $40.4\pm 0.1$.8)
RF $547.8\pm 0.6$.3)
XGB $118.8\pm 0.7$.4)
SVM $101.3\pm 1.0$.9)
DNN $574.5\pm 4.5$.8)
@@
TABLE VIII: Performance evaluation of ConCap running various port scan scenarios on a multi-node K8s cluster versus microK8s on a WSL-enabled laptop. The results show minimal differences in initialization time and resource usage between the two environments. However, the K8s cluster offers horizontal scalability, allowing for greater resource availability and enabling more parallel scenario executions.
@@
D-B Reproducing Multi-step and Multi-host attacks
Threat Model. We assume a network consisting of four or more hosts, each having different privileges and containing different information. An attacker has obtained control of one host, and wants to expand their control and steal sensitive information. However, the attacker does not know the network topology, and must hence carry out reconnaissance activities to identify vulnerabilities that can be exploited to reach the intended goal. In other words, we envision an attack chain spanning multiple stages and hosts, which can be mapped to a range MITRE ATT&CK tactics mitreattack . Altogether, the attack encompasses the following stages (expressed via the well-known MITRE ATT&CK terminology): Gather Victim Network Information, Network Service Discovery, Brute Force, Internal Proxy, SSH Lateral Movement, Data Collection from Local System, Exfiltration Over C2, and Exploitation for Privilege Escalation. Our intention is showing how to use ConCap to generate (and automatically label) NetFlows of the entire attack chain—in an end-to-end fashion.
Implementation. To replicate the setting envisioned in our threat model, we defined the scenario in Listing 5 (in the Appendix). Such a scenario embeds the use case, schematically depicted in Fig. 5, in which an attacker first performs a port scan against an SSH server using nmap nmap , then launches a brute-force SSH attack using Medusa medusa . Upon discovering valid credentials, the attacker uses SSH to set up a SOCKS proxy through the compromised host and scans additional internal hosts (a MySQL database and a vulnerable libssh server). Then, the attacker forwards internal ports via SSH tunneling to the local machine, allowing authenticated access and data exfiltration against the MySQL target and enabling the attacker to exploit a known vulnerability against libssh.
Considerations. In practice, when ConCap processes such a scenario, all of the aforementioned operations are automatically carried out. We also note that each target in this scenario is separately labeled with its function (e.g., “brute-force-ssh”, “exfiltration”, or “cve”) and configured with specific network constraints (e.g., bandwidth and delay). At the end of the entire experiment, the OpenSSH server received 3635 network packets corresponding to 184 CICFlowMeter and 185 argus NetFlows. The MySQL server received 280 packets corresponding to 109 CICFlowMeter and 110 argus NetFlows. The libSSH server received 246 packets corresponding to 106 CICFlowMeter and 106 argus NetFlows. All NetFlows (which we provide in our repository repository ) are labeled according to the specifications of the scenario. Notably, by sharing this scenario (i.e., Listing 5), future researchers can inspect the entire traffic-generation process: this is important for open science and to address the skepticism around NIDS research flood2024bad .
D-C Testing additional ML-based NIDS with ConCap
In our paper (§VI), we showed that ConCap can be used to test “simple” ML-based (e.g., RF or SVM) NIDS. Here, we expand our assessment by showing that ConCap’s generated data can be used also by other families of ML-driven NIDS.
Please note that our goal is merely to show that ConCap’s data can be used to train and test such additional methods: we do not seek to (a) propose new methods, (b) outperform/tweak existing ones, or (c) benchmark prior work.
Considered models. There are hundreds of papers proposing, or evaluating, methods to detect network intrusions via ML apruzzese2023sok . For the sake of this additional demonstration, we consider the methods considered in the recent USENIX’23 paper xNIDS wei2023xnids . Specifically, this work considers two families of deep-learning–based methods for intrusion detection: one reliant on AutoEncoders (drawn from mirsky2018kitsune ; shone2018deep ), which we denote as AE-IDS; and another one reliant on Recurrent Neural Networks (drawn from yin2017deep ; jan2020throwing ), which we denote as RNN-IDS. Note that, altogether, the papers proposing these methods have thousands of citations on Google Scholar (as of December 2025) or are published in top-tier venues (e.g., S&P’22, NDSS’18). Then, to provide yet-another perspective, we also consider transformer-based approaches which empower multimodal large-language models (LLMs), inspired by the recent luay2025multimodal .
Implementation and Datasets. We take the code from wei2023xnids (which is publicly available) as a blueprint for our AE-IDS and RNN-IDS models; whereas, for LLMs, we consider: ChatGPT 5.1, and Gemini Flash 2.5 (both of which being the best freely-available commercial models as of December 2025). For both of these LLMs, inspired by luay2025multimodal , we will assess them with a zero-shot prompting strategy (with the prompt in Listing 7), as well as in an “zero-shot-with-augmentation” fashion (with the prompt in Listing 8). [^6] Note that the approach in luay2025multimodal uses two-dimensional plots as input to the LLM, whereas we use ConCap’s raw data. This is because we want to show that ConCap can be used off-the-shelf to test LLMs. However, there is nothing preventing one from generating the same two-dimensional plots of luay2025multimodal (which simply show the source and destination IP for the x- and y-axes, and points indicate the amount of exchanged bytes—all of which being metrics derivable from ConCap’s data. For the evaluation dataset, we consider the same used in §VI-A, i.e., a mix of CICIDS17 and the data generated by ConCap referring to the ssh-patator attack. Such a setup is valid for the sake of our demonstration, given that the approaches we considered have been evaluated in completely different setups (e.g., the “outdated” NSL-KDD or the CIC-DoS2017 jazi2017detecting ).
Evaluation and Results. We replicate the experiments in §VI-A. Specifically, we train each model (AE-IDS and RNN-IDS) on 80% of the benign data in CICIDS17, as well as on 80% of the data generated by ConCap for ssh-patator (launched with default options); and then we test it on the remaining 20% (benign and malicious). We also consider doing a mixed experiment by testing each model on 20% of the ssh-patator included in CICIDS17; as well as by creating another variant trained on 80% of the ssh-patator included in CICIDS17 and testing it on 20% of the ssh-patator generated by ConCap. For the LLMs, the zero-shot prompt does not have any training (barring that to develop the LLM itself), whereas the augmented-zero-shot prompt embeds a fine-tuning step done on the same data used to train the AE-IDS and RNN-IDS. We repeat our experiments five times to ensure consistency. Let us discuss the results of our experiments:
  • AE-IDS: This experiment confirms our conclusions drawn from §VI-A. Specifically, when trained [^7] Training the AE-IDS required $\approx$10 minutes on our system. on ConCap’s generated data (which is always malicious), the AE-IDS achieves tprtpr=0.999 on the testing partition of ConCap’s generated data, and tprtpr=0.987 on the testing partition of malicious data of the same attack included in CICIDS17; conversely, when trained on the malicious data of CICIDS17, the AE-IDS achieves tprtpr=0.955 on the testing partition of CICIDS17, and tprtpr=0.952 on the testing partition of ConCap’s generated data. The fprfpr is 0.478 on the former case, and 0.209 in the latter case: such an underwhelming result is because our AE-IDS uses the same thresholding mechanism used in xNIDS wei2023xnids (note that AutoEncoders are not classifiers: to use them in a classification task, one must specify a threshold on the reconstruction error. We used the one of xNIDS, which was derived on a different dataset. To improve the fprfpr, one can simply change the threshold). The different fprfpr is because the number of datapoints in the training set is different across the two experiments, but what is crucial is that, for this detector, the data generated by ConCap is essentially equivalent to that of CICIDS17, since the classification results on these two groups are statistically the same (verified with a t-test: $p<.05$).
  • RNN-IDS: For this experiment, we report that, after 2 hours of training, the underlying RNN did not reach convergence. Therefore, given our previous experiment, as well as those in our main paper (§VI-A) showing that simple models (such as RF or SVM) can achieve near-perfect detection accuracy while requiring seconds to train (see Table VII(b)), we conclude that using RNN-IDS on our dataset is not a wise choice and stopped this experiment. However, what matters is that the RNN-IDS was being trained, meaning that ConCap’s data was “compatible” with that expected in RNN-IDS, thereby validating the goal of these experiments (note that SaTML’26 call for papers explicitly states to be mindful of the computing resources spent in training ML models, and we believe that additional training would not have changed our conclusions).
  • ChatGPT: When assessed in the “zero-shot” prompt, the model always classified each datapoint as benign. This is expected: the malicious datapoints are just repeated SSH attempts, which not necessarily imply malicious intent. However, by using the “augmented-zero-shot” prompt (which uses ConCap’s generated labelled data to “teach” the LLM that quick bursts of SSH connections are malicious), the model achieves a perfect classification accuracy (i.e., tprtpr=1.000 and fprfpr=0.000). We inspected the output provided by ChatGPT, and found that, to derive such a result, the model learned heuristics that enabled a clear separation of the benign from malicious datapoints. (note that these results align with those in luay2025multimodal ).
  • Gemini: The results for Gemini 2.5 match those of ChatGPT.
The code of these experiments is in our repository repository .
Takeaway. Our experiments show that ConCap can be used to develop and assess multiple families of ML-based NIDS. Moreover, our results further confirm that ConCap’s generated data can be functionally equivalent to that generated in other environments (e.g., that of CICIDS17).
Listing 7: Zero-shot prompt used to evaluate LLMs with ConCap’s data.
You are a cybersecurity analyst working in a Security Operations Center (SOC). You are analyzing a CSV file that contains NetFlow records.
Data Description:
Each row in the CSV represents a network flow with features such as: Src Port, Dst Port, Protocol, Flow Duration, Number of packets Your task is to classify each row as either: 0 = Benign 1 = Malicious
Your Tasks:
Determine whether the NetFlow record shows suspicious behavior. Output only 0 or 1 for each row (0 = benign, 1 = malicious) on a separate row and write to csv file.
Listing 8: Augmented zero-shot prompt used to evaluate LLMs with ConCap’s data. We augment the prompt in Listing 7 by instructing the LLM to use labeled data (in the X_train.csv file) as a guide for its inference.
You are a cybersecurity analyst working in a Security Operations Center (SOC). You are analyzing a CSV file that contains NetFlow records (X_test_shuffled.csv).
Data Description:
Each row in the CSV represents a network flow with features such as: Src Port, Dst Port, Protocol, Flow Duration, Number of packets Your task is to classify each row as either: 0 = Benign 1 = Malicious
Your Tasks:
Determine whether the NetFlow record shows suspicious behavior. Base your predictions on the provided labeled NetFlows (X_train.csv and y_train.csv). Output only 0 or 1 for each row (0 = benign, 1 = malicious) on a separate row and write to csv file.